Suppose that as an economist working for the Bureau of Labor Statistics youre given the following information for the current year: Private Disposable Income -- Federal Budget Deficit National Saving Total Domestic Investment- Total Tax Revenue Depreciation Investment Transfer Payments Gov. Interest Payments Net Factor Payments $20,000 $2,000 $3,000 $12,000 $5,000 $3,000* $1,500 S0 $1,000 *Depreciation Investment is included in total investment. With these figures, compute the following: (a)What is the level of private saving and consumption in this economy? (b)What is the level of government purchases? (c)What is the amount of net exports and the current account. Is the country running a foreign trade surplus or deficit? (d) Compute the amount of Gross Domestic Product that this economy is generating for the given year. (e) Besides GDP, we are also interested in some other measures of income. Compute GNP and Net National Product (f)Suppose that a technology company needs to finance an additional $1000 of new capital. If the current account, government saving, and private disposable income remain the same what must happen to private saving and consumption? (g) Suppose the governments budget deficit increases to $3,000. Assuming the level of Private Saving, Investment, and Net Factor Payments remain the same, calculate Net Exports. Is there a twin deficits phenomena?
Answer #1

3.a) (Y - T+TR) = private disposable income = 20,000

Or, Y = 20,000 + T - TR = 20,000 +5000 - 1500 = 23,500

Federal budget deficit = 2000, that means (T - G - TR ) = -2000. Therefore, G = T - TR + 2000 = 5000 - 1500 + 2000 = 5500.

Total Investment (I) = total domestic investment + depreciation investment = (12000+3000) = 15000

Now, we know that, National savings = (Y - C - G) = I

Therefore, C = Y - G - I = $(23,500 - 5500 - 15,000) = $3000

And, private savings = (Y - T +TR - C) = $(23,500 - 5000 + 1500 - 3000) = $17,000

Therefore, level of private savings is $17,000 and Consumption is $3000.

B) Level of Government purchases (G) = $5500 (calculated in part A).

C) For this economy, Y = $23,500, C = $3000 , G = $5500, I = $15,000.

We know that, Y = C + I + G + Net export (NX)

Therefore, NX = Y - C - G - I = $(23,500 - 3000 - 5500 -15000) = 0

It means the economy is neither running a trade surplus nor trade deficit.

D) GDP(Y)= $23,500 (calculated in part A)